Can you put these into y=mx+b
x-y=0
5x-3y=10
and x-2y=1
5x-4y= -23
and 4x+3y=8
x-2y=13
and 2x+3y= -12
8x-5y=40
please and thank you

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Just solve for y.
Get y = ---------- whatever.

Answer by stanbon(75887)

put these into y=mx+b
x-y=0
y = x
-----------------
5x-3y=10
3y = 5x-10
y = (5/3)x-(10/3)
-------------------------
and x-2y=1
2y = x-1
y = (1/2)x-(1/2)
-------------------------
5x-4y= -23
4y = 5x+23
y = (5/4)x+(23/4)
-------------------------
and 4x+3y=8
3y = -4x+8
y = (-4/3)x + (8/3)
---------------------------
x-2y=13
2y = x-13
y = (1/2)x-(13/2)
---------------------------
and 2x+3y= -12
3y = -2x - 12
y = (-2/3)x -4
-------------------------
8x-5y=40
5y = 8x - 40
y = (8/5)x-8
==================
Cheerd,
Stan H.